# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ 5.1 Case 1
For the first experiment we consider a 2-component mixture of normal distribution for random effect where mean of one component is twice the mean of the other component. So we expect that some components degrade significantly faster than others, which is also evident from Figure 2(a). For semi-parametric method we take truncation number $N$ as the number of units in the study that is 10. The mean, standard deviation (s.d) and 95% Credible interval of random effects for each of 10 units, fixed effect $\alpha$ and variance of error $\sigma_{\epsilon}^{2}$ computed based on generated posterior samples are presented in Table 1.
Table 1: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,...,10$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$)
@@
$\beta_{1}$ 3.048 0.465 (2.137, 3.981) 2.968 0.458 (2.054, 3.864)
$\beta_{2}$ 5.485 0.463 (4.571, 6.410) 5.238 0.479 (4.289, 6.198)
$\beta_{3}$ 2.309 0.466 (1.397, 3.246) 2.285 0.487 (1.345, 3.278)
$\beta_{4}$ 3.559 0.451 (2.690, 4.458) 3.453 0.452 (2.590, 4.346)
$\beta_{5}$ 4.192 0.449 (3.304, 5.075) 4.041 0.458 (3.166, 4.927)
$\beta_{6}$ 5.627 0.457 (4.763, 6.529) 5.341 0.488 (4.391, 6.293)
$\beta_{7}$ 2.955 0.465 (2.060, 3.910) 2.903 0.457 (2.016, 3.771)
$\beta_{8}$ 3.324 0.455 (2.469, 4.260) 3.216 0.454 (2.312, 4.103)
$\beta_{9}$ 3.365 0.453 (2.478, 4.263) 3.246 0.455 (2.345, 4.118)
$\beta_{10}$ 6.960 0.472 (6.033, 7.928) 6.596 0.538 (5.521, 7.616)
$\alpha$ 1.878 0.141 (1.591, 2.154) 1.956 0.154 (1.662, 2.261)
$\sigma_{\epsilon}^{2}$ 0.707 0.141 (0.481, 1.037) 0.677 0.142 (0.443, 0.986)
It is evident from Table 1, that for the proposed semi-parametric method the estimated posterior mean of fixed effect is closer to the true value compared to the parametric method. On the other side, estimated posterior mean for error variance is higher than true value for both the methods. It is observed that the units that are slowly degrading is characterized by random effects with low posterior means and significantly higher posterior means corroborate to the fact that the unit is degrading faster. Observe that for fast degrading units, posterior mean of random effect is little higher for parametric method compared to semi-parametric method. Also for the slowly degrading units the same pattern follows although the posterior mean is very close for both the methods. The posterior samples generated are used to estimate residual lifetime distribution of each individual unit. For each unit we apply the above-mentioned TMCMC technique to simulate observations from the target distribution. For each case we take $p=\frac{1}{2}$.
Figure 3: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced by semi-parametric method.
Figure 4: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced by parametric method.
Table 2: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method
Unit True value Parametric method Semi-parametric method
median 95% Predictive interval median 95% Predictive interval
Unit 1 0.665 1.362 (0.548, 2.388) 1.417 (0.514, 2.490)
Unit 2 0.023 0.211 (0.001, 0.428) 0.222 (0.001, 0.453)
Unit 3 0.894 2.094 (0.812, 3.967) 2.128 (0.789, 4.238)
Unit 4 0.710 1.012 (0.378, 1.812) 1.055 (0.385, 1.855)
Unit 5 0.607 0.700 (0.207, 1.309) 0.745 (0.248, 1.380)
Unit 6 0.024 0.175 (0.001, 0.377) 0.209 (0.001, 0.441)
Unit 7 0.886 1.457 (0.618, 2.599) 1.437 (0.520, 2.545)
Unit 8 0.917 1.158 (0.449, 2.028) 1.191 (0.407, 2.125)
Unit 9 0.551 1.122 (0.413, 1.968) 1.164 (0.521, 2.153)
Unit 10 0.025 0.158 (0.001, 0.393) 0.162 (0.001, 0.394)
The samples produced by TMCMC technique for both parametric and semi-parametric model are used for prediction. We collect the samples according to the discussion done in Section 4.2 and they are used for prediction. Predictive intervals are constructed using the 95% highest probability density region of the generated samples. The estimates for residual lifetime are rounded off to three decimal places are presented in Table 2 along with the true residual lifetime values for each unit. Figure 3 and 4 shows histograms of the generated residual lifetimes of Unit 1,2 and 3 along with the trace plot of and autocorrelation plot of generated samples for semi-parametric and parametric method respectively. Histograms and autocorrelation plots are generated by samples at lag size 10 for each unit. In each case we consider the median of the samples as the predicted residual lifetime value for each individual unit. It is evident from Table 2 that for most of the units both parametric and semi-parametric method produce intervals containing the true value.
We observe that error in predicting the residual lifetime using both the proposed semi-parametric and parametric method is nearly equal for slowly degrading units. On the other side, prediction error for parametric method is lesser than semi-parametric method for faster degrading units. Comparison of prediction error for both the methods is provided in Appendix D.
5.2 Case 2
In second case we consider a 3-component mixture of normal distribution for random effect, so we expect three different cluster of units with significant different degradation rates, which is also evident from Figure 2(b). For semi-parametric method we take truncation number $N$ as 30. The mean, standard deviation (s.d) and 95% credible interval computed based on generated posterior samples for random effects, fixed effect $\alpha$ and variance of measurement error $\sigma_{\epsilon}^{2}$ are presented in Table 3.
@@ 5.2 Case 2
Table 3: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,...,30$), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$)
@@
$\beta_{1}$ 4.374 0.453 (3.496, 5.270) 4.388 0.435 (3.553, 5.269)
$\beta_{2}$ 3.723 0.443 (2.884, 4.611) 3.808 0.450 (2.946, 4.712)
$\beta_{3}$ 2.408 0.439 (1.531, 3.250) 2.642 0.491 (1.712, 3.622)
$\beta_{4}$ 4.034 0.441 (3.146, 4.874) 4.096 0.446 (3.237, 5.001)
$\beta_{5}$ 2.811 0.441 (1.946, 3.683) 3.003 0.463 (2.145, 3.941)
$\beta_{6}$ 5.429 0.430 (4.589, 6.281) 5.330 0.468 (4.367, 6.130)
$\beta_{7}$ 3.373 0.444 (2.519, 4.242) 3.511 0.456 (2.653, 4.405)
$\beta_{8}$ 2.586 0.443 (1.744, 3.494) 2.804 0.477 (1.881, 3.805)
$\beta_{9}$ 4.886 0.460 (3.977, 5.813) 4.834 0.447 (3.941, 5.711)
$\beta_{10}$ 5.316 0.442 (4.485, 6.218) 5.240 0.458 (4.393, 6.149)
$\beta_{11}$ 2.627 0.449 (1.750, 3.510) 2.843 0.490 (1.909, 3.826)
$\beta_{12}$ 6.197 0.445 (5.351, 7.072) 6.017 0.504 (5.002, 7.011)
$\beta_{13}$ 4.743 0.434 (3.897, 5.584) 4.703 0.440 (3.849, 5.593)
$\beta_{14}$ 2.529 0.450 (1.656, 3.412) 2.781 0.486 (1.817, 3.757)
$\beta_{15}$ 4.127 0.443 (3.288, 5.014) 4.180 0.443 (3.304, 5.068)
$\beta_{16}$ 1.337 0.453 (0.465, 2.247) 1.696 0.559 (0.684, 2.861)
$\beta_{17}$ 2.320 0.445 (1.470, 3.225) 2.572 0.494 (1.632, 3.553)
$\beta_{18}$ 3.554 0.443 (2.670, 4.415) 3.674 0.449 (2.803, 4.579)
$\beta_{19}$ 4.023 0.433 (3.177, 4.887) 4.083 0.449 (3.194, 4.985)
$\beta_{20}$ 2.187 0.438 (1.315, 3.058) 2.453 0.506 (1.475, 3.470)
$\beta_{21}$ 3.408 0.437 (2.549, 4.278) 3.551 0.463 (2.667, 4.501)
$\beta_{22}$ 4.769 0.433 (3.921, 5.644) 4.755 0.458 (3.885, 5.674)
$\beta_{23}$ 2.920 0.442 (2.055, 3.799) 3.097 0.471 (2.177, 4.022)
$\beta_{24}$ 4.325 0.440 (3.492, 5.165) 4.342 0.447 (3.470, 5.237)
$\beta_{25}$ 2.191 0.439 (1.328, 3.063) 2.454 0.494 (1.499, 3.438)
$\beta_{26}$ 6.698 0.448 (5.833, 7.579) 6.469 0.518 (5.445, 7.468)
$\beta_{27}$ 5.920 0.442 (5.062, 6.786) 5.772 0.483 (4.852, 6.735)
$\beta_{28}$ 4.486 0.440 (3.639, 5.362) 4.505 0.438 (3.674, 5.404)
$\beta_{29}$ 1.553 0.446 (0.677, 2.453) 1.872 0.548 (0.834, 2.999)
$\beta_{30}$ 6.633 0.441 (5.786, 7.508) 6.391 0.516 (5.392, 7.419)
$\alpha$ 2.091 0.086 (1.922, 2.257) 2.051 0.116 (1.813, 2.268)
$\sigma_{\epsilon}^{2}$ 0.745 0.079 (0.605, 0.921) 0.761 0.074 (0.577, 1.028)
It is evident from Table 3 that for the proposed semi-parametric method the estimated posterior mean of fixed effect is closer to the true value for both parametric and semi-parametric method. The estimated posterior mean for error variance is higher than true value for both the methods, which is also evident in the Case 1 analysis. Observe that for fastest degrading units, posterior mean of random effect is higher for parametric method compared to semi-parametric method, for example see units 6 and 27. For unit 2 and 4 which can be considered as a moderate speed degrading units, it is evident that the semi-parametric produces higher posterior means of random effects compared to parametric model. The same pattern is also evident for the slow degrading units also, for example see units 20 and 25.
Figure 5: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by Semi-parametric method.
Figure 6: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by parametric method.
@@
Table 4: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method
@@
Unit True value Parametric method Semi-parametric method
median 95% Predictive interval median 95% Predictive interval
Unit 1 0.071 0.579 (0.127, 1.087) 0.487 (0.094, 0.924)
Unit 2 0.259 0.841 (0.265, 1.514) 0.675 (0.196, 1.220)
Unit 3 1.261 1.832 (0.735, 3.372) 1.626 (0.635, 2.935)
Unit 4 0.354 0.711 (0.185, 1.310) 0.688 (0.155, 1.237)
Unit 5 0.508 1.472 (0.601, 2.615) 1.089 (0.392, 1.958)
Unit 6 0.021 0.286 (0.001, 0.624) 0.321 (0.002, 0.680)
Unit 7 0.858 1.035 (0.308, 1.827) 0.978 (0.341, 1.768)
Unit 8 0.842 1.651 (0.613, 2.958) 1.454 (0.524, 2.719)
Unit 9 0.021 0.158 (0.020, 0.322) 0.166 (0.016, 0.326)
Unit 10 0.023 0.210 (0.003, 0.438) 0.226 (0.001, 0.471)
Unit 11 0.466 1.638 (0.657, 3.028) 0.959 (0.334, 1.724)
Unit 12 0.022 0.231 (0.001, 0.502) 0.267 (0.002, 0.562)
Unit 13 0.038 0.228 (0.034, 0.461) 0.241 (0.033, 0.471)
Unit 14 0.984 1.742 (0.727, 3.175) 1.511 (0.563, 2.800)
Unit 15 0.094 0.334 (0.086, 0.611) 0.335 (0.066, 0.612)
Unit 16 1.968 4.243 (1.532, 10.864) 3.166 (0.855, 7.597)
Unit 17 1.060 1.981 (0.787, 3.591) 1.684 (0.616, 3.103)
Unit 18 0.152 0.966 (0.327, 1.703) 0.348 (0.188, 0.998)
Unit 19 0.096 0.288 (0.074, 0.526) 0.282 (0.072, 0.516)
Unit 20 1.398 2.133 (0.820, 3.991) 1.827 (0.598, 3.436)
Unit 21 0.149 1.034 (0.407, 1.909) 0.596 (0.222, 1.110)
Unit 22 0.024 0.187 (0.008, 0.358) 0.193 (0.016, 0.377)
Unit 23 0.339 1.370 (0.556, 2.407) 0.780 (0.271, 1.409)
Unit 24 0.088 0.236 (0.052, 0.452) 0.243 (0.036, 0.451)
Unit 25 0.994 2.139 (0.800, 3.975) 1.813 (0.675, 3.473)
Unit 26 0.022 0.167 (0.001, 0.419) 0.199 (0.001, 0.475)
Unit 27 0.021 0.199 (0.001, 0.488) 0.226 (0.001, 0.541)
Unit 28 0.036 0.227 (0.045, 0.442) 0.226 (0.038, 0.433)
Unit 29 1.037 3.458 (1.253, 7.828) 2.721 (0.813, 6.098)
Unit 30 0.021 0.116 (0.001, 0.328) 0.148 (0.001, 0.423)
Similar to the previous analysis in Case 1, we generate observations from the residual lifetime distribution using TMCMC technique with $p=\frac{1}{2}$. Predictive intervals are constructed for each unit using the 95% highest probability density region of the generated samples. As in the previous case, we consider the median of the samples as the predicted value for each individual unit. Table 4 presents residual lifetime estimates rounded off to three decimal places along with the true values for each 30 unit. Histograms of the generated residual lifetimes of Unit 1,2 and 3 along with the trace plot and autocorrelation plot for the generated samples for semi-parametric and parametric method is presented in Figure 5 and 6 respectively. Histograms and autocorrelation plots are generated by samples at lag size 10 for each unit in this case also. We observe that in case of fast degrading units, predicted value by the parametric method is more accurate compared to the semi-parametric method for example see Unit 26,30. For the slow degrading units, it is evident that the predicted residual lifetime by the parametric method is higher than true values compared to the semi-parametric method, for example, see unit 8,11. Similar pattern is also observed for the moderate degrading units, for example see unit 1,4. But there are also some moderate degrading units for example unit 9, 13 where the predicted value for both the method is very close to each other. It is also observed that for most of the cases produced predictive intervals for both the methods contain true value. Comparison of prediction error for the both parametric and semi-parametric method is provided in Appendix D.
6 Data Analysis
Now we demonstrate analysis of fatigue crack size dataset given in Lu & Meeker[2]. The measurements are taken for 21 units. For each unit the initial crack size was 0.9 inch and data collection was terminated at the first inspection after a unit’s crack size reached 1.6 inches or censored after 0.12 million cycles, whichever came first. Degradation paths of the units are presented in Figure 7. It may be noted from the plot that some of the units are degrading faster than other units, so it can be considered that the units are coming from heterogeneous population. For semi-parametric method we take truncation number $N$ as 21.
Figure 7: Development of crack sizes over time [2]

### v2

@@ 6.1 Case 1
For the first experiment we consider a 3-component mixture of gamma distribution for random effect where mean of first component is 2, second component is 3.5 and third component is 5. So we expect that some units degrade significantly faster than others, which is also evident from Figure 1(a). Here we present the analysis for $n=10$ and $m=31$ with respect to prior 2 and 5 for semi-parametric model and prior 2 and 3 for parametric model. The mean, standard deviation (s.d) and 95% Credible interval for random effects for each of 10 units, fixed effect $\alpha$ and variance of error $\sigma_{\epsilon}^{2}$ computed based on generated posterior samples under prior 2 are presented in Table 4, whereas posterior estimates with respect to prior 3 of parametric model and prior 5 for semi-parametric model are given in Table 5. The trace plot and autocorrelation plot for the generated mcmc samples of $\alpha$, $\beta_{i}$, $i=1,2,3$ and $\sigma_{\epsilon}^{2}$ for semi-parametric and parametric method are provided in Figure 15 in Appendix B.
Table 4: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 2 for both parametric and semi-parametric method
@@
$\beta_{1}$ 2.345 0.282 (1.820, 2.888) 2.339 0.291 (1.749, 2.923)
$\beta_{2}$ 6.166 0.279 (5.628, 6.716) 6.157 0.289 (5.602, 6.764)
$\beta_{3}$ 2.461 0.288 (1.906, 2.998) 2.465 0.276 (1.941, 2.986)
$\beta_{4}$ 4.751 0.277 (4.229, 5.286) 4.748 0.279 (4.241, 5.309)
$\beta_{5}$ 3.403 0.284 (2.838, 3.982) 3.399 0.266 (2.890, 3.923)
$\beta_{6}$ 1.860 0.282 (1.309, 2.404) 1.856 0.277 (1.335, 2.445)
$\beta_{7}$ 1.993 0.271 (1.472, 2.512) 1.996 0.279 (1.452, 2.595)
$\beta_{8}$ 3.542 0.287 (2.975, 4.087) 3.531 0.274 (2.992, 4.093)
$\beta_{9}$ 3.204 0.266 (2.672, 3.730) 3.203 0.281 (2.659, 3.770)
$\beta_{10}$ 2.458 0.269 (1.925, 2.985) 2.454 0.280 (1.902, 3.000)
$\alpha$ 1.951 0.089 (1.783, 2.124) 1.952 0.09 (1.773, 2.121)
$\sigma_{\epsilon}^{2}$ 0.653 0.072 (0.527, 0.801) 0.649 0.071 (0.526, 0.797)
@@
Table 5: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 3 for parametric and prior 5 for semi-parametric method
@@
$\beta_{1}$ 2.326 0.283 (1.743, 2.917) 2.335 0.286 (1.792, 2.906)
$\beta_{2}$ 6.189 0.281 (5.639, 6.759) 6.179 0.277 (5.664, 6.752)
$\beta_{3}$ 2.453 0.277 (1.936, 2.986) 2.454 0.282 (1.927, 3.005)
$\beta_{4}$ 4.781 0.267 (4.233, 5.285) 4.773 0.281 (4.224, 5.352)
$\beta_{5}$ 3.422 0.271 (2.896, 3.964) 3.421 0.278 (2.843, 3.961)
$\beta_{6}$ 1.866 0.278 (1.311, 2.437) 1.846 0.281 (1.301, 2.443)
$\beta_{7}$ 1.981 0.269 (1.471, 2.511) 1.988 0.275 (1.453, 2.505)
$\beta_{8}$ 3.544 0.268 (3.021, 4.074) 3.556 0.261 (3.067, 4.058)
$\beta_{9}$ 3.206 0.267 (2.667, 3.730) 3.209 0.287 (2.636, 3.766)
$\beta_{10}$ 2.467 0.278 (1.903, 3.005) 2.453 0.283 (1.908, 2.993)
$\alpha$ 1.944 0.084 (1.767, 2.100) 1.946 0.085 (1.780, 2.114)
$\sigma_{\epsilon}^{2}$ 0.658 0.073 (0.533, 0.807) 0.656 0.069 (0.533, 0.798)
It is evident from Tables 4 and 5, that both parametric and proposed semi-parametric method produces credible interval for fixed effect and error variance, that contains the true value of the parameters. We also find in our analysis that as number of observations per unit is increased, the estimates are closer to true value. The units that are slowly degrading are characterized by random effects with low posterior means and significantly higher posterior means corroborate to the fact that the unit is degrading faster. We observe that for each unit, the posterior mean of random effects is very close for both the methods. It is found that the posterior estimates are not significantly different with respect to the prior choices as the number of observations per unit increases, so we consider prior 2 to compute residual lifetime distribution function for both the model. For each unit we apply the above-mentioned TMCMC technique to simulate observations from the target residual lifetime distribution.
@@
Table 6: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method with $n=10,m=31$
@@
. Unit True value Parametric Semi-parametric median 95% Predictive interval median 95% Predictive interval 1 2.175 2.866 (1.883, 3.991) 2.834 (1.788, 4.119) 2 0.355 0.473 (0.216, 0.780) 0.450 (0.150, 0.715) 3 3.156 2.664 (1.775, 3.709) 2.700 (1.698, 3.808) 4 0.375 0.899 (0.514, 1.277) 0.922 (0.528, 1.278) 5 1.260 1.673 (1.081, 2.274) 1.637 (1.110, 2.278) 6 2.917 3.859 (2.409, 5.969) 3.882 (2.343, 5.522) 7 2.992 3.521 (2.190, 5.052) 3.533 (2.288, 5.070) 8 1.033 1.558 (0.989, 2.130) 1.540 (1.026, 2.194) 9 1.502 1.850 (1.193, 2.539) 1.821 (1.148, 2.505) 10 2.513 2.686 (1.748, 3.908) 2.687 (1.741, 3.668)
Table 7: KS distance between true and approximated residual lifetime distribution for Bayesian parametric method and proposed Bayesian semi-parametric method
. Unit $m=11$ $m=31$ Parametric Semi-parametric Parametric Semi-parametric 1 0.2257 0.2671 0.4121 0.4217 2 0.3654 0.3389 0.3944 0.4621 3 0.3395 0.3145 0.8358 0.8239 4 0.2820 0.2741 0.2592 0.2996 5 0.4021 0.3912 0.1565 0.2107 6 0.6803 0.6706 0.3096 0.2949 7 0.7745 0.8118 0.5214 0.5324 8 0.5524 0.5154 0.0686 0.0822 9 0.3663 0.3789 0.4403 0.4678 10 0.2559 0.2325 0.6409 0.6283
Table 8: Prediction accuracy for Parametric and Semi-Parametric method
. Scenario $M_{1}$ $M_{2}$ Parametric Semi-parametric Parametric Semi-parametric RMSE MAE RMSE MAE RMSE MAE RMSE MAE $n=10,m=11$ 1.710 1.028 1.831 1.059 1.710 1.028 1.762 1.045 $n=10,m=31$ 0.527 0.690 0.520 0.681 0.536 0.693 0.541 0.698 $n=30,m=11$ 2.448 1.174 2.653 1.191 2.447 1.173 2.490 1.174 $n=30,m=31$ 1.051 0.918 1.056 0.921 1.052 0.922 1.053 0.920 $n=50,m=11$ 0.796 0.747 1.761 1.155 0.772 0.749 1.778 1.159 $n=50,m=31$ 1.040 0.899 1.036 0.898 1.033 0.897 1.046 0.901
We generate samples according to the discussion done in Section 4.2 and they are used for prediction. Predictive intervals are constructed using the 95% highest probability density region of the generated samples. The histograms of the generated residual lifetimes for units 1-3 along with the trace plot of and autocorrelation plot of generated samples for semi-parametric and parametric method are provided in Figure 3 and 4 in the Appendix A to this article. It is evident from Table 6 that for most of the units both parametric and semi-parametric method produce intervals containing the true value, although predictive interval produced for unit 4 does not contain true value for both the parametric and semi-parametric method. The reason for this is both unimodal and mixture of normal distributions for random effect is unable to model gamma mixtures properly.
Kolmogorov-Smirnov (KS) distance obtained for $m=11,31$ and $n=10$ is presented in Table 7. We observe that for the slowly degrading units for example unit 3, 6 semi-parametric method produced smaller KS value compared to parametric method. The reason for this is that the mixture of normal is able to estimate the associated random effect generated from the mixing component with smallest mean, whereas the unimodal distribution for parametric method is unable to estimate this. On the other side, KS distance for parametric method is lesser for faster degrading units, due to the fact that the mean of the posterior distribution of random effects is shifted towards the higher value for this method and hence it produces more accurate estimates for the units coming from the component with highest mean. We evaluate overall prediction accuracy using RMSE and MAE for each combination of $n$ and $m$. It is evident from Table 8 that, as the number of observations per unit increases for each $n=10,30$ and 50, both RMSE and MAE decreases under $M_{1}$ and $M_{2}$ for most of the cases. Both the proposed semi-parametric and parametric method for $M_{1}$ and $M_{2}$, is nearly equal as both the observations and sample size increases.
6.2 Case 2
For the Case 2, We consider a 3-component mixture of Weibull distribution for random effect. The degradation observation generated for this case is given in Figure 1(b) for $n=10,m=31$. The mean, standard deviation (s.d) and 95% Credible interval of random effects for each of 10 units, fixed effect $\alpha$ and variance of error $\sigma_{\epsilon}^{2}$ computed based on generated posterior samples with respect to the same prior combinations as discussed for Case 1 are presented in Tables 9 and 10 respectively. The trace plot and autocorrelation plot for the generated mcmc samples of $\alpha$, $\beta_{i}$, $i=1,2,3$ and $\sigma_{\epsilon}^{2}$ for semi-parametric and parametric method are provided in Figure 16 in Appendix B.